OpenStudy (anonymous):

The sum of 2 interior angles of this triangle is 140°. What is the measure of the third interior angle? A. 30° B. 70° C. 50° D. 40°

OpenStudy (anonymous):

D

sam (.sam.):

The sum of all interior angles of a triangle is 180 degree,a triangle has 3 angles, so, 180 - 2sides = 40 degrees

OpenStudy (anonymous):

suppose,XYZ is a triangle.if x+y=140,then, x+y+z=180[the sum of all interior angles of a triangle is 180 degrees] 140+z=180[x+y=140] z=40 degrees.